### 25 Dec 2023 v1

# Finite almost simple groups whose holomorph contains a solvable regular subgroup

Abstract.
In our previous paper, we gave a complete list of the finite non-abelian simple groups whose holomorph contains a solvable regular subgroup. In this paper, we refine our previous work by considering all finite almost simple groups. In particular, our result yields a complete characterization of the finite almost simple groups which occur as the type of a Hopf-Galois structure on a solvable extension, or equivalently, the additive group of a skew brace with solvable multiplicative group.
1. Introduction
Let $G$ and $N$ be two finite groups of the same order. We say that the pair $(G,N)$ is realizable if the holomorph of $N$ contains a regular subgroup which is isomorphic to $G$. Recall that the holomorph of $N$ is defined to be
$\mathrm{Hol}(N)=\rho(N)\rtimes\mathrm{Aut}(N)=\lambda(N)\rtimes\mathrm{Aut}(N)$
as a subgroup of the symmetric group of $N$, where $\lambda$ and $\rho$ are the left and right regular representations of $N$. By [5, 2] and [6], respectively, it is known that $(G,N)$ is realizable if and only if there exists a Hopf-Galois structure of type $N$ on a Galois $G$-extension of fields, or equivalently, a skew brace with additive group $N$ and multiplicative group isomorphic to $G$. This is exactly our motivation of studying realizability (see [4] and [7] for more details).
For a pair $(G,N)$ to be realizable, it is conjectured that if $G$ is insolvable then $N$ is also insolvable (see [8, 3] for some progress on this). On the other hand, the converse is known to be false – there exist realizable pairs $(G,N)$ for which $G$ is solvable but $N$ is insolvable. It seems to be a natural problem to try and determine the insolvable groups $N$ that can occur in such a pair. In a previous paper [9, Theorems 1.2 & 1.3], we proved that:
Theorem 1.1.
Let $N$ be a finite almost simple group. If $(G,N)$ is realizable for some solvable group $G$, then the socle of $N$ is isomorphic to
  1. $(1)$
    $\mathrm{PSL}_{3}(3),\,\mathrm{PSL}_{3}(4),\,\mathrm{PSL}_{3}(8),\,\mathrm{PSU}_{3}(8),\,\mathrm{PSU}_{4}(2),\,\mathrm{M}_{11}$; or
  2. $(2)$
    $\mathrm{PSL}_{2}(q)$ with $q\neq 2,3$ a prime power.
The converse is also true when $N$ is a finite non-abelian simple group.
The purpose of this paper is to study the converse of Theorem 1.1 for an arbitrary finite almost simple group $N$ that is not necessarily simple. As it turns out, the proposition below (see [9, Corollary 2.2 & Proposition 2.7]) is enough to do the job except when $N=\mathrm{PSU}_{3}(8)$. This case has already been dealt with in [9] and so may be ignored.
Proposition 1.2.
Let $N$ be any finite group.
  1. $(a)$
    If $(G,N)$ is realizable for a solvable group $G$, then there exists a subgroup
    $\mathrm{Inn}(N)\leq P\leq\mathrm{Aut}(N)$
    such that $P=AB$ is factorized by some solvable subgroups $A,\,B$ which satisfy $A\mathrm{Inn}(N)=B\mathrm{Inn}(N)$.
  2. $(b)$
    If $N$ has trivial center and there exists a subgroup
    $\mathrm{Inn}(N)\leq P\leq\mathrm{Aut}(N)$
    such that $P=AB$ is factorized by some solvable subgroups $A,\,B$ which satisfy $A\cap B=1,\,A\mathrm{Inn}(N)=B\mathrm{Inn}(N)$, and $A$ splits over $A\cap\mathrm{Inn}(N)$, then $(G,N)$ is realizable for some solvable group $G$.
Here $\mathrm{Inn}(N)$ denotes the inner automorphism group of $N$.
The converse of Theorem 1.1 does not hold in general, but we have a very nice criterion and here is our main result.
Theorem 1.3.
Let $N$ be a finite almost simple group whose socle $\mathrm{Soc}(N)$ is isomorphic to one of the non-abelian simple groups in Theorem 1.1. There is a solvable group $G$ such that $(G,N)$ is realizable if and only if
  1. $(1)$
    $\mathrm{Soc}(N)\simeq\mathrm{PSL}_{3}(3),\,\mathrm{PSL}_{3}(8),\,\mathrm{PSU}_{4}(2),\,\mathrm{M}_{11}$; or
  2. $(2)$
    $\mathrm{Soc}(N)\simeq\mathrm{PSL}_{2}(q)$ with $q\neq 2,3$ a prime power; or
  3. $(3)$
    $\mathrm{Soc}(N)\simeq\mathrm{PSL}_{3}(4)$ with $[N:\mathrm{Soc}(N)]\neq 2,4$; or
  4. $(4)$
    $\mathrm{Soc}(N)\simeq\mathrm{PSL}_{3}(4)$ with $[N:\mathrm{Soc}(N)]=2$ and $N\unlhd\mathrm{Aut}(\mathrm{Soc}(N))$; or
  5. $(5)$
    $\mathrm{Soc}(N)\simeq\mathrm{PSU}_{3}(8)$ with $[N:\mathrm{Soc}(N)]\neq 2,6$; or
  6. $(6)$
    $\mathrm{Soc}(N)\simeq\mathrm{PSU}_{3}(8)$ with $[N:\mathrm{Soc}(N)]=6$ and $N\unlhd\mathrm{Aut}(\mathrm{Soc}(N))$.
Remark 1.4.
In the language of Hopf-Galois structures and skew braces, our result yields a complete classification of the finite almost simple groups that can occur as the type of a Hopf-Galois structure on a solvable extension, or equivalently, the additive group of a skew brace with solvable multiplicative group.
By definition, an almost simple group $N$ corresponds to a subgroup of the outer automorphism group of its socle $\mathrm{Soc}(N)$. Let us note that
$\mathrm{Out}(\mathrm{PSL}_{3}(3))\simeq C_{2},\,\ \mathrm{Out}(\mathrm{PSL}_{3}(4))\simeq D_{12},\,\ \mathrm{Out}(\mathrm{PSL}_{3}(8))\simeq C_{6},$
$\mathrm{Out}(\mathrm{PSU}_{3}(8))\simeq C_{3}\times S_{3},\,\ \mathrm{Out}(\mathrm{PSU}_{4}(2))\simeq C_{2},\,\ \mathrm{Out}(\mathrm{M}_{11})\simeq C_{1}.$
Moreover, for a prime power $q=p^{f}$ with $q\neq 2,3$, we have
$\mathrm{Out}(\mathrm{PSL}_{2}(q))\simeq C_{(2,q-1)}\times C_{f}.$
Here, as usual $C_{n}$ is the cyclic group of order $n$, $D_{2n}$ is the dihedral group of order $2n$, and $S_{n}$ is the symmetric group on $n$ letters. (See [10] for example.)
Therefore, the converse of Theorem 1.1 holds with only five exceptions of $N$ up to isomorphism. The exceptions are the following:
  • The subgroups $\mathrm{PSL}_{3}(4)\leq N\leq\mathrm{Aut}(\mathrm{PSL}_{3}(4))$ corresponding to
    $\langle s\rangle,\,\ \langle rs\rangle,\,\ \langle r^{3},s\rangle.$
    Note that $\langle r^{3}\rangle$ is also a subgroup of $D_{12}$ of order $2$, but this is normal
    and so falls under case (4) of Theorem 1.3. Here $r$ and $s$ are the usual generators of $D_{12}$ with $r^{6}=s^{2}=1$ and $srs=r^{-1}$.
  • The subgroups $\mathrm{PSU}_{3}(8)\leq N\leq\mathrm{Aut}(\mathrm{PSU}_{3}(8))$ corresponding to
    $\{1\}\times\langle(1\,2)\rangle\mbox{ and }C_{3}\times\langle(1\,2)\rangle.$
    Note that $\{1\}\times S_{3}$ is also a subgroup of $C_{3}\times S_{3}$ of order $6$, but this is normal and so falls under case (6) of Theorem 1.3.
We now move on and use Proposition 1.2 to prove Theorem 1.3. We use the help of Magma [1] except for the infinite family $\mathrm{PSL}_{2}(q)$ with $q\neq 2,3$.
2. Proof for $\mathrm{PSL}_{3}(3),\,\mathrm{PSL}_{3}(4),\,\mathrm{PSL}_{3}(8),\,\mathrm{PSU}_{3}(8),\,\mathrm{PSU}_{4}(2),\,\mathrm{M}_{11}$
We ran the following code in Magma [1].
T:=(the non-abelian simple group in question);
AutT:=PermutationGroup(AutomorphismGroup(T));
InnT:=Socle(AutT);
NN:=[N‘subgroup:N in Subgroups(AutT)|InnT subset N‘subgroup];
//A list of the finite almost simple groups N with socle T
//up to conjugacy class in Aut(T).
for N in NN do
  AutN:=Normalizer(AutT,N);
  //We identify Aut(N) with the normalizer of N in Aut(T)
  //and in particular Inn(N) with N.
  PP:=[P‘subgroup:P in Subgroups(AutN)|N subset P‘subgroup];
  //A list of the subgroups P between Inn(N) and Aut(N)
  //up to conjugacy class in Aut(N).
  PPa:=[]; PPb:=[];
  //A list to contain the P which satisfies the conditions
  //in Propositions 1.2(a) and 1.2(b), respectively.
  for P in PP do
    SolSub:=[S‘subgroup:S in SolvableSubgroups(P)];
    if exists{<A,B>:A,B in SolSub|
         #P eq #A*#B/#(A meet B) and
         sub<P|A,N> eq sub<P|B,N>} then
      Append(~PPa,P);
    end if;
    if exists{<A,B>:A,B in SolSub|
         #(A meet B) eq 1 and
         #P eq #A*#B and
         sub<P|A,N> eq sub<P|B,N> and
         HasComplement(A,A meet N)} then
      Append(~PPb,P);
    end if;
  end for;
  if IsEmpty(PPa) then
    conclusion:="false";
    //Proposition 1.2(a): If PPa is empty then (G,N) is not
    //realizable for any solvable group G.
    //Thus the converse of Theorem 1.1 is "false" for N.
  elif not IsEmpty(PPb) then
    conclusion:="true";
    //Proposition 1.2(b): If PPb is non-empty then (G,N) is
    //realizable for some solvable group G.
    //Thus the converse of Theorem 1.1 is "true" for N.
  else
    conclusion:="inconclusive";
    //Proposition 1.2 is not sufficient to determine if the
    //converse of Theorem 1.1 holds for N.
  end if;
  if IsNormal(AutT,N) then
    normalORnot:="normal";
  else
    normalORnot:="not normal";
  end if;
 <#N/#InnT,normalORnot,conclusion>;
end for;
For $T=\mathrm{PSL}_{2}(3),\,\mathrm{PSL}_{3}(8),\,\mathrm{PSU}_{4}(2),\,\mathrm{M}_{11}$, the output of conclusion is all true and this yields case (1) of Theorem 1.3.
For $T=\mathrm{PSL}_{3}(4)$ and $[N:\mathrm{Soc}(N)]\neq 2,4$, the output of conclusion is all true and this yields case (3) of Theorem 1.3.
For $T=\mathrm{PSL}_{3}(4)$ and $[N:\mathrm{Soc}(N)]=2,4$, the output is as follows:
<2, "normal", "true">
<2, "not normal", "false">
<2, "not normal", "false">
<4, "not normal", "false">
We then deduce case (4) of Theorem 1.3.
For $T=\mathrm{PSU}_{3}(8)$ and $[N:\mathrm{Soc}(N)]\neq 1,2,6$, the output of conclusion is all true and this yields case (5) of Theorem 1.3 except when $N\simeq\mathrm{PSU}_{3}(8)$, which is already covered by Theorem 1.1 so this case may be ignored. Let us remark that the output of conclusion is inconclusive in this case and [9] dealt with it using a slightly different method.
For $T=\mathrm{PSU}_{3}(8)$ and $[N:\mathrm{Soc}(N)]=2,6$, the output is as follows:
<2, "not normal", "false">
<6, "normal", "true">
<6, "not normal", "false">
We then deduce case (6) of Theorem 1.3.
This proves cases (1), (3), (4), (5), (6) of Theorem 1.3.
3. Proof for $\mathrm{PSL}_{2}(q)$ with $q\neq 2,3$ a prime power
Let $q=p^{f}$ be a prime power with $q\neq 2,3$ and let
$\mathrm{PSL}_{2}(q)\leq N\leq\mathrm{Aut}(\mathrm{PSL}_{2}(q)).$
Since $\mathrm{Out}(\mathrm{PSL}_{2}(q))$ is abelian, such an $N$ is normal in $\mathrm{Aut}(\mathrm{PSL}_{2}(q))$, so we may identify $\mathrm{Aut}(N)$ with $\mathrm{Aut}(\mathrm{PSL}_{2}(q))$, and in particular $\mathrm{Inn}(N)$ with $N$. By Proposition 1.2(b), it is then enough to show that there is a subgroup
$N\leq P\leq\mathrm{Aut}(\mathrm{PSL}_{2}(q))$
for which there exist solvable subgroups $A,\,B$ of $P$ satisfying
(3.1) $P=AB,\,\ A\cap B=1,\,\ AN=BN,\,\ \mbox{and $A$ splits over $A\cap N$}.$
Note that the last two conditions are trivial when $P$ is taken to be $N$ since then $AN=N=BN$ and $A$ obviously splits over $A\cap N=A$.
We may regard $\mathrm{PSL}_{2}(q)$ as a normal subgroup of $\mathrm{PGL}_{2}(q)$ via the natural embedding, and in particular $\mathrm{PGL}_{2}(q)$ as a subgroup of $\mathrm{Aut}(\mathrm{PSL}_{2}(q))$ which acts on $\mathrm{PSL}_{2}(q)$ via conjugation. It is well-known that
(3.2) $\mathrm{Aut}(\mathrm{PSL}_{2}(q))=\mathrm{PGL}_{2}(q)\rtimes F,$
where $F\simeq C_{f}$ is the subgroup consisting of the automorphisms induced by the field automorphisms of $\mathbb{F}_{p^{f}}$ over $\mathbb{F}_{p}$. By the proof of [9, Theorem 1.3], we already know that there are solvable subgroups $C,\,D$ of $\mathrm{PGL}_{2}(q)$ such that
$\mathrm{PGL}_{2}(q)=CD,\,\ C\cap D=1,\,\ C\cdot\mathrm{PSL}_{2}(q)=D\cdot\mathrm{PSL}_{2}(q).$
Explicitly, we may take $C=\widetilde{C}/Z$ and $D=\widetilde{D}/Z$, where
$\widetilde{C}$ =$=\left\{\begin{bmatrix}x&-dy\\ y&x-cy\end{bmatrix}:x,y\in\mathbb{F}_{q},\,(x,y)\neq(0,0)\right\},$
$\widetilde{D}$ =$=\left\{\begin{bmatrix}u&v\\ 0&w\end{bmatrix}:u,w\in\mathbb{F}_{q}^{\times},\,v\in\mathbb{F}_{q}\right\}.$
Here $Z$ denotes the center of $\mathrm{GL}_{2}(q)$ and $X^{2}+cX+d$ is the minimal polynomial over $\mathbb{F}_{q}$ of some generator of $\mathbb{F}_{q^{2}}^{\times}$. Observe that
$\theta\cdot\begin{bmatrix}x&y\\ z&w\end{bmatrix}Z\cdot\theta^{-1}=\begin{bmatrix}\theta(x)&\theta(y)\\ \theta(z)&\theta(w)\end{bmatrix}Z$
in the group (3.2) for all $\theta\in F$ and $\left[\begin{smallmatrix}x&y\\ z&w\end{smallmatrix}\right]\in\mathrm{GL}_{2}(q)$. It in particular implies that $D$ is normalized by $F$. Thus, we have that $DE=D\rtimes E$ is a subgroup of $\mathrm{Aut}(\mathrm{PSL}_{2}(q))$ for any subgroup $E$ of $F$. In what follows, let us take $E$ to be the projection of $N$ onto $F$ along $\mathrm{PGL}_{2}(q)$, and let
$P=\mathrm{PGL}_{2}(q)\rtimes E.$
Clearly $N$ is contained $P$. For the choices of $A,\,B$, we separate into cases.
First suppose that $q\equiv 0\pmod{4}$. Then $\mathrm{PGL}_{2}(q)=\mathrm{PSL}_{2}(q)$ and so
$N=\mathrm{PGL}_{2}(q)\rtimes E=P.$
Take $(A,B)=(C,D\rtimes E)$. That $A,\,B$ are solvable, $P=AB$, and $A\cap B=1$ are all clear. The last two conditions in (3.1) are trivial here since $P=N$.
Next suppose that $q\equiv 1,3\pmod{4}$. It was shown in [9, Section 4] that
(3.3) $\begin{cases}C\mbox{ splits over }C\cap\mathrm{PSL}_{2}(q)&\mbox{when }q\equiv 1\pmod{4},\\ D\mbox{ splits over }D\cap\mathrm{PSL}_{2}(q)&\mbox{when }q\equiv 3\pmod{4}.\end{cases}$
Here we shall take
$(A,B)=\begin{cases}(C,D\rtimes E)&\mbox{when }q\equiv 1\pmod{4},\\ (D\rtimes E,C)&\mbox{when }q\equiv 3\pmod{4}.\end{cases}$
That $A,\,B$ are solvable, $P=AB$, and $A\cap B=1$ are all obvious. Since
$C\cdot\mathrm{PSL}_{2}(q)=D\cdot\mathrm{PSL}_{2}(q)\mbox{ and }\mathrm{PSL}_{2}(q)\leq N,$
we clearly have $CN=DN$. Note that $\mathrm{PGL}_{2}(q)=\mathrm{PSL}_{2}(q)D$ holds because $D$ contains an element lying outside of $\mathrm{PSL}_{2}(q)$, for example
$\begin{bmatrix}u&0\\ 0&1\end{bmatrix}Z,\mbox{ where }u\in\mathbb{F}_{q}^{\times}\setminus(\mathbb{F}_{q}^{\times})^{2}.$
By definition, elements of $E$ may be represented as $\pi\eta$ for $\pi\in\mathrm{PGL}_{2}(q)$ and $\eta\in N$. Since $N$ contains $\mathrm{PSL}_{2}(q)$, we then deduce that $E$ lies in
$\mathrm{PGL}_{2}(q)N=\mathrm{PSL}_{2}(q)DN=DN.$
Thus, we have $(D\rtimes E)N=DN=CN$ and this yields $AN=BN$. Finally, we check that $A$ indeed splits over $A\cap N$.
  • $q\equiv 1\pmod{4}:$ Here $A=C$. Since $A\leq\mathrm{PGL}_{2}(q)$, we have
    $A\cap N=\begin{cases}A\cap\mathrm{PSL}_{2}(q)&\mbox{if }N\cap\mathrm{PGL}_{2}(q)=\mathrm{PSL}_{2}(q),\\ A&\mbox{if }N\cap\mathrm{PGL}_{2}(q)=\mathrm{PGL}_{2}(q).\end{cases}$
    In the former case $A$ splits over $A\cap N$ by (3.3), and in the latter case $A$ certainly splits over $A\cap N=A$.
  • $q\equiv 3\pmod{4}:$ Here $A=D\rtimes E$. Since $q=p^{f}$, the exponent $f$ must be odd in this case, and $\mathrm{Out}(\mathrm{PSL}_{2}(q))\simeq C_{2}\times C_{f}\simeq C_{2f}$ is cyclic. We then see that there are only two possibilities for $N$, namely
    $N=\mathrm{PSL}_{2}(q)\rtimes E\mbox{ or }N=\mathrm{PGL}_{2}(q)\rtimes E.$
    In the latter case $N=P$ and $A$ certainly splits over $A\cap N=A$. In the former case, we observe that
    $A\cap N$ $=(D\rtimes E)\cap(\mathrm{PSL}_{2}(q)\rtimes E)=(D\cap\mathrm{PSL}_{2}(q))\rtimes E$
    and in particular
    $[A:A\cap N]$ $=[D:D\cap\mathrm{PSL}_{2}(q)].$
    Here, instead of applying (3.3) directly, we shall need to use its proof given in [9, Section 4]. It was shown by noting that
    $[D:D\cap\mathrm{PSL}_{2}(q)]=2,\,\ |D\cap\mathrm{PSL}_{2}(q)|=\frac{q(q-1)}{2}$
    and by using the Schur-Zassenhaus theorem. But then we have
    $[A:A\cap N]=2,\,\ |A\cap N|=\frac{q(q-1)}{2}\cdot|E|.$
    Since $|E|$, being a divisor of $f$, is odd in this case, we can similarly use the Schur-Zassenhaus theorem to deduce that $A$ splits over $A\cap N$.
In both cases, we have shown that $A$ splits over $A\cap N$ for our choice of $A$.
We have thus verified the existence of $P$ satisfying (3.1) in all cases. This proves case (2) of Theorem 1.3.

### 21 Mar 2024 v2

@@
 In this paper, we refine our previous work by considering all finite almost simple groups.
-In particular, our result yields a complete characterization of the finite almost simple groups which occur as the type of a Hopf-Galois structure on a solvable extension, or equivalently, the additive group of a skew brace with solvable multiplicative group.
+In particular, our result yields a complete characterization of the finite almost simple groups which occur as the type of a Hopf-Galois structure on a solvable extension, or equivalently, the additive group of a skew brace having a solvable multiplicative group.
 1. Introduction
 Let $G$ and $N$ be two finite groups of the same order.
-We say that the pair $(G,N)$ is realizable if the holomorph of $N$ contains a regular subgroup which is isomorphic to $G$.
+We say that the pair $(G,N)$ is realizable if the holomorph of $N$ contains a regular subgroup that is isomorphic to $G$.
 Recall that the holomorph of $N$ is defined to be
@@
 This is exactly our motivation of studying realizability (see [4] and [7] for more details).
-For a pair $(G,N)$ to be realizable, it is conjectured that if $G$ is insolvable then $N$ is also insolvable (see [8, 3] for some progress on this).
+For the pair $(G,N)$ to be realizable, it is conjectured that if $G$ is insolvable then $N$ is also insolvable (see [8, 3] for some progress on this).
 On the other hand, the converse is known to be false – there exist realizable pairs $(G,N)$ for which $G$ is solvable but $N$ is insolvable.
@@
     $\mathrm{Inn}(N)\leq P\leq\mathrm{Aut}(N)$
-    such that $P=AB$ is factorized by some solvable subgroups $A,\,B$ which satisfy $A\mathrm{Inn}(N)=B\mathrm{Inn}(N)$.
+    such that $P=AB$ is factorized by some solvable subgroups $A$ and $B$ that satisfy the equality $A\mathrm{Inn}(N)=B\mathrm{Inn}(N)$.
   2. $(b)$
@@
     $\mathrm{Inn}(N)\leq P\leq\mathrm{Aut}(N)$
-    such that $P=AB$ is factorized by some solvable subgroups $A,\,B$ which satisfy $A\cap B=1,\,A\mathrm{Inn}(N)=B\mathrm{Inn}(N)$, and $A$ splits over $A\cap\mathrm{Inn}(N)$, then $(G,N)$ is realizable for some solvable group $G$.
+    such that $P=AB$ is factorized by some solvable subgroups $A$ and $B$ that satisfy $A\cap B=1,\,A\mathrm{Inn}(N)=B\mathrm{Inn}(N)$, and $A$ splits over $A\cap\mathrm{Inn}(N)$, then $(G,N)$ is realizable for some solvable group $G$.
 Here $\mathrm{Inn}(N)$ denotes the inner automorphism group of $N$.
@@
   1. $(1)$
+    $\mathrm{Soc}(N)\simeq\mathrm{PSL}_{2}(q)$ with $q\neq 2,3$ a prime power; or
+  2. $(2)$
     $\mathrm{Soc}(N)\simeq\mathrm{PSL}_{3}(3),\,\mathrm{PSL}_{3}(8),\,\mathrm{PSU}_{4}(2),\,\mathrm{M}_{11}$; or
-  2. $(2)$
-    $\mathrm{Soc}(N)\simeq\mathrm{PSL}_{2}(q)$ with $q\neq 2,3$ a prime power; or
   3. $(3)$
@@
 Moreover, for a prime power $q=p^{f}$ with $q\neq 2,3$, we have
-$\mathrm{Out}(\mathrm{PSL}_{2}(q))\simeq C_{(2,q-1)}\times C_{f}.$
+$\mathrm{Out}(\mathrm{PSL}_{2}(q))\simeq C_{\gcd(2,q-1)}\times C_{f}.$
 Here, as usual $C_{n}$ is the cyclic group of order $n$, $D_{2n}$ is the dihedral group of order $2n$, and $S_{n}$ is the symmetric group on $n$ letters. (See [10] for example.)
-Therefore, the converse of Theorem 1.1 holds with only five exceptions of $N$ up to isomorphism. The exceptions are the following:
+Therefore, the converse of Theorem 1.1 holds with only five exceptions of $N$, as follows.
+Here, only the isomorphism class of $N$ matters, so it suffices to consider the subgroups of $\mathrm{Out}(\mathrm{Soc}(N))$ up to conjugacy.
   • The subgroups $\mathrm{PSL}_{3}(4)\leq N\leq\mathrm{Aut}(\mathrm{PSL}_{3}(4))$ corresponding to
@@
 We now move on and use Proposition 1.2 to prove Theorem 1.3. We use the help of Magma [1] except for the infinite family $\mathrm{PSL}_{2}(q)$ with $q\neq 2,3$.
-2. Proof for $\mathrm{PSL}_{3}(3),\,\mathrm{PSL}_{3}(4),\,\mathrm{PSL}_{3}(8),\,\mathrm{PSU}_{3}(8),\,\mathrm{PSU}_{4}(2),\,\mathrm{M}_{11}$
+2. Proof of cases (2) to (6)
 We ran the following code in Magma [1].
@@
   PPa:=[]; PPb:=[];
-  //A list to contain the P which satisfies the conditions
+  //A list to contain the P that satisfies the conditions
   //in Propositions 1.2(a) and 1.2(b), respectively.
@@
 We then deduce case (6) of Theorem 1.3.
-This proves cases (1), (3), (4), (5), (6) of Theorem 1.3.
-3. Proof for $\mathrm{PSL}_{2}(q)$ with $q\neq 2,3$ a prime power
+3. Proof of case (1)
 Let $q=p^{f}$ be a prime power with $q\neq 2,3$ and let
@@
 Note that the last two conditions are trivial when $P$ is taken to be $N$ since then $AN=N=BN$ and $A$ obviously splits over $A\cap N=A$.
-We may regard $\mathrm{PSL}_{2}(q)$ as a normal subgroup of $\mathrm{PGL}_{2}(q)$ via the natural embedding, and in particular $\mathrm{PGL}_{2}(q)$ as a subgroup of $\mathrm{Aut}(\mathrm{PSL}_{2}(q))$ which acts on $\mathrm{PSL}_{2}(q)$ via conjugation.
+We may regard $\mathrm{PSL}_{2}(q)$ as a normal subgroup of $\mathrm{PGL}_{2}(q)$ via the natural embedding, and in particular $\mathrm{PGL}_{2}(q)$ as a subgroup of $\mathrm{Aut}(\mathrm{PSL}_{2}(q))$ that acts on $\mathrm{PSL}_{2}(q)$ via conjugation.
 It is well-known that
@@
 $P=\mathrm{PGL}_{2}(q)\rtimes E.$
-Clearly $N$ is contained $P$. For the choices of $A,\,B$, we separate into cases.
+Clearly $P$ contains $N$. For the choices of $A,\,B$, we separate into cases.
 First suppose that $q\equiv 0\pmod{4}$. Then $\mathrm{PGL}_{2}(q)=\mathrm{PSL}_{2}(q)$ and so
@@
     In the former case $A$ splits over $A\cap N$ by (3.3), and in the latter case $A$ certainly splits over $A\cap N=A$.
-  • $q\equiv 3\pmod{4}:$ Here $A=D\rtimes E$.
-Since $q=p^{f}$, the exponent $f$ must be odd in this case, and $\mathrm{Out}(\mathrm{PSL}_{2}(q))\simeq C_{2}\times C_{f}\simeq C_{2f}$ is cyclic.
-We then see that there are only two possibilities for $N$, namely
+  • $q\equiv 3\pmod{4}:$ Here $A=D\rtimes E$. Since $q=p^{f}$, the exponent $f$ must be odd in this case, and so
+    $\mathrm{Out}(\mathrm{PSL}_{2}(q))\simeq C_{2}\times C_{f}\simeq C_{2f}$
+    is cyclic. We then see that there are only two possibilities for $N$, namely
     $N=\mathrm{PSL}_{2}(q)\rtimes E\mbox{ or }N=\mathrm{PGL}_{2}(q)\rtimes E.$